OpenStudy (anonymous):

not finished but just want to make sure im doing this right? Write the equation of the line passing through the points (-6,1) and (-4,2). m = 2 - 1/ -4 - (-6) = 1/ 2 = 0.5

OpenStudy (anonymous):

m = 2 - 1/ -4 - (-6) = 1/ 2 = 0.5 is this correct ?

OpenStudy (anonymous):

is it linear?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you mean slope above?

OpenStudy (anonymous):

im plugged the numbers into the formula m = y2 - y1 / x2 - x1

OpenStudy (anonymous):

am i right? y=1/2x+4

OpenStudy (anonymous):

no